OpenStudy (matimaticas):

Find the product of (x + (3+5i))^2

OpenStudy (anonymous):

simply use this... (a+b)^2=a^2+b^2+2ab

OpenStudy (anonymous):

and for i^2=-1 just plug it if you get i^2

OpenStudy (matimaticas):

thx!!

OpenStudy (anonymous):

you welcoem...